OpenStudy (anonymous):

7n+6=4n-9 please show work.

OpenStudy (anonymous):

you want to get n, so you solve it for n: 7n+6=4n-9 bring all n's on one side and all numbers on the other: 7n-4n= -9-6 3n= -15 n=-5 so you get to n=-5